OpenStudy (anonymous):

lim(xto0)2[(e^x-(1+x))/x^2]=

OpenStudy (anonymous):

\[\lim_{x\to 0} e^x-\frac{(1+x)}{x^2}\] ?

OpenStudy (anonymous):

satellite u lost 2

OpenStudy (anonymous):

ooh i know \[lim_{x\to 0} \frac{2e^x-(1+x)}{x^2}\]

OpenStudy (anonymous):

i bet that is it. l'hopital's rule problem. of the form \[\frac{0}{0}\]

OpenStudy (anonymous):

take derivative top and bottom separate. you will need to do it twice because after the first time you still get \[\frac{0}{0}\]

myininaya (myininaya):

the top is 2e^0-1-0=2(1)-1-0=2-1=1

OpenStudy (anonymous):

first time \[\frac{2e^x-1}{2x}\]

OpenStudy (anonymous):

still \[\frac{0}{0}\]

OpenStudy (anonymous):

second time \[\frac{2e^x}{2}\] let x ->0 get 1

OpenStudy (anonymous):

oops i am totally wrong. sorry. first time it is \[\frac{1}{0}\] no limit. so sorry.

OpenStudy (anonymous):

must be my bed time.

OpenStudy (anonymous):

yes you have lost the derivative of constant 1 first time

OpenStudy (anonymous):

by first time i mean before you start. it is not \[\frac{0}{0}\] to begin with so no l'hopital. sorry

OpenStudy (anonymous):

WHERE IS THE DERIVATIVE OF CONSTANT?

OpenStudy (anonymous):

Mmmm....looking carefully at the question as posted, we are looking for the limit of 2(e^x - 1 - x)/x^2. As x goes to zero we get 2(e^0 - 1 - 0)/0^2 = 2(1 - 1)/0 = 0/0, so use l'Hopital twice to get Lim as x goes to 0 of e^x, which is 1.

myininaya (myininaya):

did you try graphing it? both sides of 0 f(x) goes to positive infinity

OpenStudy (anonymous):

WHY WE TAKE DERIVATIVE ANY REASON????